OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)?

OpenStudy (ghazi):

(x+7)(x-2)= \[x^2+7x-2x-14= x^2+5x-14\]

OpenStudy (ghazi):

use law of distribution \[(x*x+7*x-2*x-7*2)\] you'll have the above equation